# Some remarks on real algebraic maps which are topologically special generic maps

### v4

@@ Abstract.
We consider the zero set $\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
The last part of STEP 1 implies that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function. Around a point of the form $(x,y)\in\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way.
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value. We explain about this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then for some variable $x_{j_{0}}$, we have the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials. Remember that $f_{{\partial}_{j}}(x)$ is not zero for any $j$ except $j=i_{0}$. The value of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ for $x_{j_{0}}$ is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class. The product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$ is not zero by the property of the zero sets of the polynomials.
We apply the implicit function theorem to see that
$M_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$
is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$. Remember that this is also the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $\partial{\bar{N}}_{0}\times[0,1]\subset N(\partial{\bar{N}}_{0})$. More precisely, we can have one and consider identification in the following way thanks to arguments related to ”Definition 14” [20].
  • $\partial{\bar{N}}_{0}\times\{0\}$ is identified with $\partial N(\partial{\bar{N}}_{0})\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
  • $\partial{\bar{N}}_{0}\times\{t_{0}\}\subset\partial{\bar{N}}_{0}\times(0,1]$ is identified with the zero set of the polynomial $f_{j}-a_{0}t_{0}$ for a sufficiently small number $a_{0}>0$.
Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function: some $y_{j}$ is regarded as the variable of the target of the function.
We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin. This is a diffeomorphism. We map each point of the form $(x,0)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a diffeomorphism from ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ onto ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ and by the restriction another diffeomorphism $({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto ${\rm Int}\ {\bar{N}}_{0}\times S^{m-n}\subset{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
@@
We argue STEPs 1–3 again in a generalized way in the proof of Main Theorem 3. By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}^{n}}$ enjoying all desired properties in Main Theorems 1 and 2.
@@
Suppose a situation as in Main Theorem 1 is given.
Let $F_{S^{m-n}}:{\mathbb{R}}^{m-n+2}={\mathbb{R}}^{m-n+1}\times\mathbb{R}\rightarrow\mathbb{R}$ be a function defined canonically by a real polynomial enjoying the following. We naturally define $F_{S^{m-n},t}:{\mathbb{R}}^{m-n+1}\rightarrow\mathbb{R}$ as the restriction to ${\mathbb{R}}^{m-n+1}\times\{t\}$ for $t\in\mathbb{R}$ where ${\mathbb{R}}^{m-n+1}\times\{t\}$ and ${\mathbb{R}}^{m-n+1}$ are identified by the map mapping $(y,t)$ to $t$.
@@
  • ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the empty set for $t<0$.
  • For $a_{1},a_{2}\in\mathbb{R}$ such that $a_{1}<a_{2}$, define $(a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}<x\leq a_{2}\}$. We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)$ mapping ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively. ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
  • For any real number $\epsilon>0$, we have a positive real number $\delta>0$ and for $0<t<\delta$ and each point $p$ of ${F_{S^{m-n},t}}^{-1}(0)$, $0<||p||<\epsilon$.
@@
for a sufficiently large $T>0$.
As a more specific case, we can also have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
@@
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$. This is also a bundle which may not be smooth and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$.
@@
    Under the presented structure group, the bundle of (1) here which may not be a smooth bundle is in fact trivial.
  3. (3)
    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps.
@@
      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is homeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$. Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4,5$. Note also that for $k\neq 4$ this is diffeomorphic to $D^{k}$ if its boundary is diffeomorphic to $S^{k-1}$.
@@
      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and $D^{m-n+1}$: here $D^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
@@
We review several known facts on topologies and differentiable structures of spheres and disks shortly. See [34] and see also [12, 46, 47] for example. A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example. It is well-known that smooth manifolds homeomorphic to $S^{k}$ are classified via abstract and sophisticated theory for $k\geq 7$. If a smooth manifold which is homeomorphic to $S^{k}$ is smoothly embedded into ${\mathbb{R}}^{k+1}$ for $k\geq 1$ with $k\neq 4$, then this is diffeomorphic to $S^{k}$. A smooth manifold ”which is homeomorphic to $S^{4}$ and not diffeomorphic to $S^{4}$” is still undiscovered. A smooth manifold which is homeomorphic to $D^{k}$ and whose boundary is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ unless $k=4$. A smooth manifold which is homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ unless $k=4,5$. Remember that ”${\Delta}^{k}$ which is homeomorphic to $D^{k}$ and is not diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
@@
We can argue as STEP 1 of the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than $a$. We choose another positive number $T_{0}<a$.
We present an argument similar to one of STEP 2 of the proof of Main Theorems 1 and 2. Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
This mainly comes from the following assumption: ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
Around a point of the form $(x,y)\in(\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 of the proof of Main Theorems 1 and 2. The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not zero from the assumption, and the value in the case of Main Theorems 1 and 2. The resulting value is not zero. We can argue similarly to prove that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
We present an argument similar to one of STEP 3 of the proof of Main Theorems 1 and 2. For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$ and regard this as $\partial{\bar{N}}_{0}\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $\partial{\bar{N}}_{0}\times\{0\}$ are identified by the correspondence $x\mapsto(x,0)$.
@@
First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$.
@@
We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
@@
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
We have a homeomorphism onto a smooth manifold $M$. We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$. The resulting map is also special generic by the construction. It is also product-organized.
We define $W_{0}$ as
$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$
and this can be also defined as

### v5

@@ Abstract.
We consider the zero set $S_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
The last part of STEP 1 implies the fact that $S_{0}=\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$. We apply the implicit function theorem to see that this set is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$.
Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
Around a point of the form $(x,y)\in(\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way.
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a value which is not 00. We discuss this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ is not 00 for some variable $x_{j_{0}}$, since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary. Remember that this is also obtained via an isotopy in the $C^{1}$ class. For the variable $x_{j_{0}}$, the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials. Remember that $f_{{\partial}_{j}}(x)$ is not 00 for any $j$ except $j=i_{0}$. The value of the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is not 00.
We have checked that the set $S_{0}$ is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$. Put $M_{0}:=S_{0}$. Our desired map $f_{0}$ is defined as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $(\partial{\bar{N}}_{0})\times[0,1]\subset N(\partial{\bar{N}}_{0})$. More precisely, we can have one and consider our identification in the following way thanks to arguments related to ”Discussion 14” in [21].
  • The manifold $(\partial{\bar{N}}_{0})\times\{0\}$ is identified with $(\partial N(\partial{\bar{N}}_{0}))\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
  • The manifold $(\partial{\bar{N}}_{0})\times\{t_{0}\}\subset(\partial{\bar{N}}_{0})\times(0,1]$ is identified with the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}})-a_{0}t_{0}$ for a sufficiently small number $a_{0}>0$. We have ${\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=at_{0}={\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ at each point $(x,y)\in(((\partial{\bar{N}}_{0})\times\{t_{0}\})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}\subset({\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ ($y=(y_{1},\cdots,y_{m-n+1})$).
This means that around $N(\partial{\bar{N}}_{0})$, $f_{0}$ is a special generic map and represented as the product map of a Morse function for a natural height of an ($m-n+1$)-dimensional disk and the identity map on $\partial{\bar{N}}_{0}$ (for suitable local coordinates).
Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function. Some $y_{j}$ is regarded as the variable of the target of the function here.
We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin. This is a diffeomorphism. We map each point of the form $(x,0)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a diffeomorphism from $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ onto $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$. By the restriction, we also have another diffeomorphism $(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto $({\rm Int}\ {\bar{N}}_{0})\times S^{m-n}\subset({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$.
@@
We argue STEPs 1–3 again in a generalized way in the proof of Main Theorem 3.
By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}^{n}}$ enjoying all desired properties in Main Theorems 1 and 2.
@@
We show Main Theorem 3. This is a kind of generalizations of Main Theorem 2.
Before this, we review several known facts on topologies and differentiable structures of spheres and disks shortly. See [35] and see also [13, 47, 48] for example. A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example. It is well-known that smooth manifolds homeomorphic to $S^{k}$ are classified via abstract and sophisticated theory for $k\geq 7$. If a smooth manifold which is homeomorphic to $S^{k}$ is smoothly embedded into ${\mathbb{R}}^{k+1}$ for $k\geq 1$ with $k\neq 4$, then this is diffeomorphic to $S^{k}$. A smooth manifold ”which is homeomorphic to $S^{4}$ and not diffeomorphic to $S^{4}$” is still undiscovered. A smooth manifold which is homeomorphic to $D^{k}$ and whose boundary is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ unless $k=4$. A smooth manifold which is homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ unless $k=4,5$. Remember that ”${\Delta}^{k}$ which is homeomorphic to $D^{k}$ and is not diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
@@
Suppose that a situation same as one of Main Theorem 1 is given.
Let $F_{S^{m-n}}:{\mathbb{R}}^{m-n+2}={\mathbb{R}}^{m-n+1}\times\mathbb{R}\rightarrow\mathbb{R}$ be a function defined canonically by a real polynomial enjoying the following. We naturally define $F_{S^{m-n},t}:{\mathbb{R}}^{m-n+1}\rightarrow\mathbb{R}$ as the restriction to ${\mathbb{R}}^{m-n+1}\times\{t\}$ for $t\in\mathbb{R}$ where ${\mathbb{R}}^{m-n+1}\times\{t\}$ and ${\mathbb{R}}^{m-n+1}$ are identified by the map mapping $(y,t)$ to $y$.
  1. (1)
@@
  4. (4)
    We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively. The zero set ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ of the polynomial is non-singular and it does not contain the origin $0\in{\mathbb{R}}^{m-n+1}$. It is also homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
@@
for a sufficiently large $T>0$ and real polynomials $f_{{\partial}_{j}}$ in our proof of Main Theorems 1 and 2.
Explicitly, we can also have the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
@@
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is the composition of the restriction of $q_{f_{0}}$ to the preimage of a suitable small collar neighborhood $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$. Furthermore, the smooth bundle is a trivial smooth bundle.
@@
    The non-singular real algebraic manifold $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps. Here ${\Delta}^{k}$ is for a $k$-dimensional smooth manifold homeomorphic to $D^{k}$ ($k\geq 1$). This is diffeomorphic to $D^{k}$ for $k\neq 4,5$. This is also diffeomorphic to $D^{k}$ for $k\neq 4$ if the boundary is diffeomorphic to $S^{k-1}$.
@@
      Prepare the product $\bar{N}\times{\Delta}^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$.
@@
      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times{\Delta}^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times{\Delta}^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and ${\Delta}^{m-n+1}$: here ${\Delta}^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
@@
Hereafter, for a real number $t$ and $x\in{\mathbb{R}}^{k}$, we use the notation respecting the structure of the canonical vector space for ${\mathbb{R}}^{k}$ such as $tx\in{\mathbb{R}^{k}}$. The vector $tx$ is also a point of ${\mathbb{R}}^{k}$ of course.
@@
We can argue as STEP 1 in the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than the given positive number $a>0$. We choose another positive number $T_{0}<a$. We define $M_{0}$ as presented in the statement of Main Theorem 3.
We also define the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ as the composition of the canonical embedding to ${\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}$ as in Main Theorems 1 and 2 and their proof.
We present arguments similar to ones of STEP 2 in the proof of Main Theorems 1 and 2. Here we also respect the assumptions (1)–(4) here.
First we also have the relation
$M_{0}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
by our definitions similarly. We show that $M_{0}$ is also a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function. We discuss this more precisely. We consider the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some variable $y_{j}$. According to the property (4), ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$. It is assumed in (2) that $F_{0,S^{m-n}}(t)>0$ for $t>0$. We can know that the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ is not 00 for some variable $y_{j}$.
Around a point of the form $(x,y)\in((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 in the proof of Main Theorems 1 and 2. The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not 00 from the assumption (2), $\frac{1}{T}$, and the value in the case of Main Theorems 1 and 2. The value in the case of Main Theorems 1 and 2 means the value of the partial derivative of ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ for the variable $x_{j_{0}}$. The resulting value is not 00.
We can argue similarly to see that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
We present an argument similar to one in STEP 3 in the proof of Main Theorems 1 and 2. For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$. For example, we can consider this as the collar neighborhood of STEP 3 in the proof. We regard this as $(\partial{\bar{N}}_{0})\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $(\partial{\bar{N}}_{0})\times\{0\}$ are identified by the correspondence identifying $x$ with $(x,0)$.
@@
First we map $(\partial{\bar{N}}_{0})\times\{0\}$ to $(\partial{\bar{N}}_{0})\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$. We use the natural identification mapping $(x,0)$ to $(x,0)$ here. According to the property (3) for example, we have $((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}=(\partial{\bar{N}}_{0})\times\{0\}$.
@@
We can map each point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
@@
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
Thus we have a homeomorphism from the smooth manifold $M_{0}$ onto a smooth manifold $M$. The restriction of this to the set $M_{0}-(({\partial\bar{N_{0}}})\times\{0\})$ gives a diffeomorphism.
We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$. The resulting map is also special generic by the construction. It is also product-organized. We define a special generic map $f$ as this map. An essentially same argument also appears in Main Theorems 1 and 2 and their proof.
We define a new set $W_{0}\subset{\mathbb{R}}^{m+1}$ as
$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$
and by our assumption, this can be also defined as